# Jasper and Stella: distillation of SOTA embedding models

### v1

@@ Abstract
A crucial component of many deep learning applications (such as FAQ and RAG) is dense retrieval, in which embedding models are used to convert raw text to numerical vectors and then get the most similar text by MIPS (Maximum Inner Product Search). Some text embedding benchmarks (e.g. MTEB Muennighoff et al. (2022), BEIR Thakur et al. (2021), and AIR-Bench Chen et al. (2024)) have been established to evaluate embedding models accurately. Thanks to these benchmarks, we can use SOTA models; however, the deployment and application of these models in industry were hampered by their large vector dimensions and numerous parameters. To alleviate this problem, 1) we present a distillation technique that can enable a smaller student model to achieve good performance. 2) Inspired by MRLKusupati et al. (2024) we present a training approach of reducing the vector dimensions based on its own vectors or its teacher vectors. 3) We do simple yet effective alignment training between images and text to make our model a multimodal encoder. We trained Stella and Jasper models using the technologies above and achieved high scores on the MTEB leaderboard. We release the model and data at Hugging Face Hub [^1] https://huggingface.co/infgrad/jasper_en_vision_language_v1  [^2] https://huggingface.co/datasets/infgrad/jasper_text_distill_dataset, the training codes will be in this project  [^3] https://github.com/NLPJCL/RAG-Retrieval and the training logs are at Weights & Biases [^4] https://api.wandb.ai/links/dunnzhang0/z8jqoqpb.
@@ 1 Introduction
With the rapid development of natural language processing technologies, text embedding models play a crucial role in text representation, information retrieval, and generation tasks. By mapping words, sentences, or documents into a high-dimensional continuous space, these models enable similar texts to have closer vector representations, thus not only enhancing the manipulability of textual data but also significantly improving the performance of various downstream tasks. Especially in retrieval-enhanced generation (RAG) techniques, the ability of the embedding model directly affects the quality of the generated results.
Retrieval systems that index passages by embedded models can efficiently deploy and utilize them to achieve fast retrieval of relevant passages, as shown by Maximum Inner Product Search (MIPS). The growing interest in text embedding models in academia and industry has led to the recent release of many new models, such as E5 Wang et al. (2022)Wang et al. (2023)Wang et al. (2024), GTE Li et al. (2023), and Jina Günther et al. (2023). To provide a consistent method for comparing the accuracy of existing text embedding models several benchmarks have been created, and public leaderboards have been made available for different tasks on the HuggingFace platform, such as MTEB Muennighoff et al. (2022), BEIR Thakur et al. (2021), and AIR-Bench Chen et al. (2024). These efforts have further advanced the use of text embedding models in Natural Language Processing (NLP) tasks, contributing to the rapid development of a wide range of applications.
However, models that have excellent performance usually contain a large number of parameters and high dimensions; for example, NV-Embed-v2 Lee et al. (2024) Moreira et al. (2024), bge-en-icl Xiao et al. (2023)Li et al. (2024), and e5-mistral-7b-instruct Wang et al. (2022)Wang et al. (2023)Wang et al. (2024) have 7B parameters, and their vectors are 4096d. These features lead to slow inference and retrieval speeds. Actually, this is a trade-off between accuracy and speed.
In this paper, we use knowledge distillation Hinton et al. (2015) to let a student model learn a teacher model’s vectors by several well-designed losses. In order to further improve the student model’s performance, we use serval excellent embedding models as teacher models (i.e. concatenate multiple teacher vectors at dimension -1). This method does not need any supervised data, and the MTEB leaderboard results and in-house test dataset results show that our student models are far better than those in the same parameter. Because of the distillation of serval teacher models, our student models have large vector dimensions (i.e., the summary of teacher models’s vector dimensions). To solve this problem, we introduce a training method that can reduce the model’s dimensions. Just like distillation, this method also does not need any supervised data. This method does not even need the teacher’s vectors, just doing a self-distillation can still achieve a good performance.
We also train a siglip model to align its visual embedding with the text encoder’s token embeddings; as a result, our model can both encode images and text.
@@ 2 Methods
In this section, we describe our model architecture designs and four-stage training method.
@@ 2.1 Definitions
To be able to better introduce our model and training method, we make the following definitions:
  • student model: the text embedding model to be trained
  • teacher model: the SOTA embedding model used to teach student models to generate its vectors; this model will not be trained
  • vectors: also called text embedding or text representation, is a pooling of last hidden states. Its shape is (batch_size, vector_dimensions)
  • student vectors: student model’s vectors
  • teacher vectors: teacher model’s vectors
  • sv: abbreviation for student vectors
  • tv: abbreviation for teacher vectors
  • mse: abbreviation for Mean Square Error
  • matmul: matrix product of two tensors
  • vectors.T: View of the transposed vectors
@@ 2.2 Model Architecture
Our model architecture follows the simple and standard design of combining a language model with a vision encoder. As shown in Figure 1, it consists of four components:
@@
    A ViT image encoder that independently maps images into vision token embedding;
@@
    A pool that projects the vision tokens to the language model’s input dimension and reduces their count;
@@
    A transform encoder or decoder (e.g. BERT, GPT2, XLMRoberta, generally called a text embedding model);
@@
    Several fully connected layers (FC) that project the vectors to the specific dimension;
Figure 1: The model architecture of Jasper model
2.3 Stage1&2: Distillation from Teacher Vectors
The two stages’s target is to let student model to learn serval teacher models’s vectors. To achieve this goal, we designed three loss functions.
The first loss function is cosine loss which is formulated as follows:
[…]
The cosine_loss was designed with the simple intention of making the angle between the student vectors and teacher vectors in the higher-dimensional space as small as possible.
The second loss function is a supplement for cosine_loss. Apparently, the cosine loss value will not go down to zero, so there will always be an angle between student vectors and teacher vectors, we accept this angle but hope the similarity values for any text pair computed by student vectors and teacher vectors are the same. We call this loss function similarity_loss which is formulated as follows:
$\begin{split}\text{similarity\_loss}=mse(matmul(sv,sv.T),\\ matmul(tv,tv.T))\end{split}$ (2)
The third loss function is a supplement for cosine_loss and similarity_loss, it further reduces the requirement for student vectors. This loss function is a standard triplet_loss which is formulated as follows:
$\begin{split}\text{triplet\_loss}=\frac{1}{N}\sum_{\text{a, p, n}\in\text{triplet\_corpus}}\max(0,\\ \text{score}(\text{a},\text{p})-\text{s}(\text{a},\text{n})+\text{margin})\end{split}$ (3)
In this equation, $a$, $p$ and $n$ separately represent anchor text, positive text and negative text; $score(a,p)$ represents the cosine similarity between anchor text and positive text computed by student vectors. We use teacher vector to compute cosine similarity and then get anchor text, positive text and negative text according to the size of the cosine similarity. We generate triplet_corpus from a batch data, if the batch size is bsz, then the number of triplet_corpus (i.e. N) is:
$\text{N}=C_{C_{bsz}^{2}}^{2}$ (4)
Note that, $N$ is very large and proportional to the fourth power of bsz, so if bsz is too large, you may encounter OOM errors. In a short word, we use teacher vectors to generate many triplet data and then train student model by using triplet loss.
@@ 2.3 Stage1&2: Distillation from Teacher Vectors
The final loss is a weighted sum of those three losses. Algorithm 1 provides the pseudo-code of this distillation training and loss’s weight. The biggest advantage of distillation vectors is that we do not need any supervised data. No consideration of resource constraints, we can use trillions of unsupervised texts to do distillation training to get extreme performance for a given model size.
@@
The difference between stage1 and stage2 is the trained parameters. The stage1 only train the fully connected layer and the stage2 only train the fully connected layer and last three layers of student model.
Algorithm 1 Pseudocode of Stage1&2 in a PyTorch-like style.
def get_score_diff(vectors):
scores = torch.matmul(vectors, vectors.T)
scores = scores[torch.triu(torch.ones_like(scores), diagonal=1).bool()]
score_diff = scores.reshape((1, -1)) - scores.reshape((-1, 1))
score_diff = score_diff[torch.triu(torch.ones_like(score_diff), diagonal=1).bool()]
return score_diff
# start train
for batch in active_dataloader:
# teacher_vectors = torch.cat( stella_vectors, nv_v2_vectors )
teacher_vectors = batch.pop("teacher_vectors")
model_output = model(**batch)
# in stage1 and stage2, we only train the fc that convert the last hidden states to 12288d
## this fc is for learning teacher vectors, and other fc is for reducing dimensions
student_vectors = model_output["vectors_12288"].float()
student_vectors = F.normalize(student_vectors, p=2, dim=-1)
# cosine loss, cosine_loss_scale is 10.0 in our codes
cosine_loss = (1 - (student_vectors * teacher_vectors).sum(axis=1).mean()) * cosine_loss_scale
# similarity loss, similarity_loss_scale is 200.0 in our codes
similarity_loss = F.mse_loss(
input=torch.matmul(student_vectors, student_vectors.T),
target=torch.matmul(teacher_vectors, teacher_vectors.T),
) * similarity_loss_scale
# triplet_loss, triplet_loss_scale is 20.0 and triplet_margin is 0.015 in our codes
triplet_label = torch.where(get_score_diff(teacher_vectors) < 0, 1, -1)
triplet_loss = F.relu(get_score_diff(student_vectors) * triplet_label + triplet_margin).mean() * triplet_loss_scale
loss = cosine_loss + similarity_loss + triplet_loss
# do backward, step and zero_grad...
@@ 2.4 Stage3: Dimension Reduction
In the distillation of stage1 and stage2, we use a fully connected layer to project the student vectors to the teacher vectors’ dimension. Specifically, in Jasper model, we use stella_en_1.5B_v5 and NV-Embed-v2 Lee et al. (2024) as teacher model, their dimensions is 4096 and 8192, so the student vector dimensions is 12288 (4096+8192) which is too large. According to the Johnson–Lindenstrauss lemma Ghojogh et al. (2021), we can make the dimensions lower without compromising performance. Our method is simple, we add several fully connected layers to reduce student vector dimension. For example, if we add a fully connected layer with a shape of (hidden_size, 512), we can get vectors of 512d.
In stage3, we use similarity_loss and triplet_loss as loss functions. The reduced vector dimensions are not the same with teacher vectors, so we skip the cosine_loss. In this stage, we train all parameters. Note that, to ensure the accuracy of the last 2 stage vectors (i.e., vectors of 12288d), this vector still be trained with all three loss functions.
Besides the above approach of dimension reduction, we can also consider vectors of the student model as teacher vectors (a sort of self-distillation). By using this approach, we might get a little bit of performance degradation, but we can reduce the dimensionality of any embedding model just by using unsupervised data and itself. Because this paper mainly introduces the training method of the Stella and Jasper model, so we didn’t do experiments to argue the merits of the method.
2.5 Stage4: Visualized Jasper Model
In this stage, we use image caption data as training data and only train the visual encoder. The essence of this training is still distillation: the vector of caption is teacher vector, and the vector of image is student vector. The loss function is the same with stage1 (i.e. cosine_loss, similarity_loss and triplet_loss).
As in stage3, we have multiple fully connected layers, we have multiple student vectors. During training, we compute the loss for each student vector and teacher vector, and then the average of multiple losses is the final loss. The pseudo-code can be found in Algorithm 2. We think there is a lot of room for improvement in this approach, which will be explained in the Discussion section.
Algorithm 2 Pseudocode of Stage4 in a PyTorch-like style.
def get_score_diff(vectors):
scores = torch.matmul(vectors, vectors.T)
scores = scores[torch.triu(torch.ones_like(scores), diagonal=1).bool()]
score_diff = scores.reshape((1, -1)) - scores.reshape((-1, 1))
score_diff = score_diff[torch.triu(torch.ones_like(score_diff), diagonal=1).bool()]
return score_diff
# start train
for batch in active_dataloader:
# get teacher_vectors
with torch.no_grad():
teacher_vectors = model.encode(batch.pop("image_captions"))
# get all_student_vectors
all_student_vectors = model(**batch)["all_vectors"]
# for each vector we compute loss
loss = 0.0
for student_vectors in all_student_vectors:
# cosine loss, cosine_loss_scale is 10.0 in our codes
cosine_loss = (1 - (student_vectors * teacher_vectors).sum(axis=1).mean()) * cosine_loss_scale
# similarity loss, similarity_loss_scale is 200.0 in our codes
similarity_loss = F.mse_loss(
input=torch.matmul(student_vectors, student_vectors.T),
target=torch.matmul(teacher_vectors, teacher_vectors.T),
) * similarity_loss_scale
# triplet_loss, triplet_loss_scale is 20.0 and triplet_margin is 0.015 in our codes
triplet_label = torch.where(get_score_diff(teacher_vectors) < 0, 1, -1)
triplet_loss = F.relu(get_score_diff(student_vectors) * triplet_label + triplet_margin).mean() * triplet_loss_scale
loss += (cosine_loss + similarity_loss + triplet_loss)
loss /= len(all_student_vectors)
# do backward, step and zero_grad...

### v2

@@ Abstract
A crucial component in many deep learning applications, such as Frequently Asked Questions (FAQ) and Retrieval-Augmented Generation (RAG), is dense retrieval. In this process, embedding models transform raw text into numerical vectors. However, the embedding models that currently excel on text embedding benchmarks, like the Massive Text Embedding Benchmark (MTEB), often have numerous parameters and high vector dimensionality. This poses challenges for their application in real-world scenarios. To address this issue, we propose a novel multi-stage distillation framework that enables a smaller student embedding model to distill multiple larger teacher embedding models through three carefully designed losses. Meanwhile, we utilize Matryoshka Representation Learning (MRL) to reduce the vector dimensionality of the student embedding model effectively. Our student model named Jasper with 2 billion parameters, built upon the Stella embedding model, obtained the No.3 position on the MTEB leaderboard (as of December 24, 2024), achieving average 71.54 score across 56 datasets. We have released the model and data on the Hugging Face Hub  [^1] https://huggingface.co/infgrad/jasper_en_vision_language_v1  [^2] https://huggingface.co/datasets/infgrad/jasper_text_distill_dataset, and the training codes are available in this project repository  [^3] https://github.com/NLPJCL/RAG-Retrieval.
@@ 1 Introduction
With the rapid development of natural language processing technologies, text embedding models play a crucial role in text representation Kashyap et al. (2024), information retrieval Zhao et al. (2024a), and text generation tasks Gao et al. (2023). By mapping words, sentences, or documents into a high-dimensional continuous space, these models bring similar texts closer together in their vector representations, thereby not only enhancing the manipulability of textual data but also significantly improving the performance of various downstream tasks Agarwal et al. (2024); Wang et al. (2024); Zhou et al. (2024).
However, embedding models that demonstrate excellent performance on the METB leaderboard [^4] https://huggingface.co/spaces/mteb/leaderboard Muennighoff et al. (2023) usually contain a large number of parameters and high vector dimensions. For instance, both NV-Embed-v2 Lee et al. (2024); Moreira et al. (2024) and bge-en-icl Xiao et al. (2023); Li et al. (2024) have 7 billion parameters and 4096-dimensional vector representations. These characteristics lead to slow inference speeds and high storage costs, posing a significant challenge to their direct practical application.
To address the aforementioned challenges, we propose a novel multi-stage knowledge distillation framework for embedding models. Knowledge distillation is widely recognized for enhancing the effectiveness of dense retrieval training Hofstätter et al. (2021); Lin et al. (2021). In our framework, we introduce three carefully designed loss functions to distill knowledge from the teacher model to the student model. These loss functions shift from a specific constraint to a broader constraint. The first, cosine loss, calculates the absolute difference in text representations between the student and teacher models. The pointwise signal derived from a single text is straightforward, yet its limited optimization direction tends to readily lead to overfitting on the training data. Thus, we introduce the similarity loss, which measures the semantic discrepancies between the student and teacher models from a text-pair perspective. Additionally, we design the relative similarity distillation loss to further leverage relative ranking information. This ensures that the student model learns the teacher’s ranking preferences across all potential positive and negative text pairs within the batch, thereby improving the robustness of embedding learning.
To further improve the performance of the student model, we utilize multiple powerful large embedding models as teachers. Specifically, we concatenate the vectors produced by all teacher models to create the final ground truth, which inevitably leads to an increase in the student model’s vector dimension. To address this issue, we adopt a Matryoshka Representation Learning (MRL)-based training method Kusupati et al. (2024) to effectively compress the student model’s vector representation. Additionally, to develop the multimodal retrieval capability of our student model, we integrate a vision encoder and introduce a self-distillation mechanism to align the visual embeddings with the textual embeddings. In terms of the overall training process, we employ a 4-stage distillation approach to progressively transfer knowledge from the teacher models to the student model. Each stage focuses on specific aspects, combining three loss functions and fine-tuning different parameters of the student model to ensure a smooth and effective distillation process.
Experimental results on the MTEB leaderboard demonstrate that our student model named Jasper with 2 billion (2B) parameters, primarily built upon the foundation of the Stella embedding model, delivers excellent performance (average 71.54 score across 56 datasets) comparable to other embedding models with 7 billion (7B) parameters, and significantly outperforms models with fewer than 2B parameters.
The main contributions of this paper can be summarized as follows:
  • (1)
    We propose a novel multi-stage distillation framework, which enables a smaller student embedding model to effectively distill knowledge from multiple larger teacher embedding models through three carefully designed loss functions.
  • (2)
    Our 2B Jasper model obtained the No.3 position on the MTEB leaderboard (as of December 24, 2024), producing results comparable to other top-ranked 7B embedding models and significantly outperforming other models with less than 2B parameters.
@@ 2.1 Definitions
For a more comprehensive introduction of our model and distillation framework, we make the following definitions:
  • Student Model: The text embedding model that is the subject of training, tasked with learning to produce effective vector representations.
  • Teacher Model: The state-of-the-art (SOTA) embedding model serving as a teacher, guiding the student model in generating effective vectors. Notably, the teacher model will not be trained.
  • $s_{x}$: The normalized vector representation of a text $x$ produced by the student model.
  • $t_{x}$: The vector representation of the same text $x$, first normalized, then concatenated, and normalized again, produced by multiple teacher models.
  • $S_{X}$: A matrix of normalized vector representations for a batch of text $X$ produced by the student model.
  • $T_{X}$: A corresponding matrix of vector representations for the same batch of text $X$, first normalized, then concatenated, and subsequently normalized again, generated by multiple teacher models.
@@ 2.2 Model Architecture
Our student model architecture follows the simple and standard design of combining a language model with a vision encoder. As shown in Figure 1, it consists of four components:
@@
    A encoder-based language model that generates text embeddings through mean pooling.
@@
    A vision encoder that independently maps images into vision token embeddings.
@@
    A pooler that maps vision token embeddings to the same dimension as the language model’s input textual embeddings, while reducing the length of visual token sequences.
@@
    Several fully connected (FC) layers that project the embeddings to a specific dimension for the final output.
Figure 1: The model architecture of Jasper model.
2.3 Stage 1&2: Distillation from Multiple Teachers
In the first two stages of distillation, we use a fully connected layer to map the vectors of the student model onto the dimensions of the teacher models. Specifically, we employ NV-Embed-v2555https://huggingface.co/nvidia/NV-Embed-v2 and stella_en_1.5B_v5666https://huggingface.co/dunzhang/stella_en_1.5B_v5 as teacher models, which have vector dimensions of 4096 and 8192, respectively. After the mapping process, the student model’s vector dimension is adjusted to 12288, equal to the combined vector dimensions of two teacher models (4096 + 8192).
The objective of the first two stages is to enable the student model to effectively learn text representations from multiple teacher models by aligning its output vectors with the corresponding teacher vectors. To achieve this goal, we carefully design three loss functions that progress from a specific to a broader perspective. The first loss function is cosine loss, which is formulated as follows:
$\mathcal{L}_{cosine}=\sum_{x}1-s_{x}\cdot t_{x}.$ (1)
The $\mathcal{L}_{\text{cosine}}$ is designed to minimize the angular difference between student and teacher vectors in the high-dimensional space, with the aim of aligning their absolute text representations. However, the $\mathcal{L}_{cosine}$ value generally does not converge to zero, suggesting a persistent angular discrepancy between the student and the teachers. Meanwhile, the pointwise signal derived from a single text has a limited optimization direction, which can easily lead to overfitting on the training data.
$\mathcal{L}_{sim}=MSE(S_{X}S_{X}^{T},T_{X}T_{X}^{T}))$ (2)
To complement the limitations of $\mathcal{L}_{cosine}$, we introduce the second loss function, similarity loss, as defined in (2), which models the semantic matching differences between the student and teacher models from a text-pair perspective. This loss function ensures a relatively consistent judgment of similarity between the student model and the teacher models, without enforcing an absolute fit between the student model and the teacher model.
$\begin{split}\mathcal{L}_{resim}=\frac{1}{N}\sum_{t_{i}\cdot t_{j}>t_{m}\cdot t_{n}}MAX(0,\\ s_{m}\cdot s_{n}-s_{i}\cdot s_{j}+margin)\end{split}$ (3)
To further leverage relative comparison signals, inspired by CoSENT loss [^7] https://spaces.ac.cn/archives/8847, we propose the third loss function, relative similarity distillation loss, as defined in (3). For each batch of text data, we employ teacher models to automatically generate soft labels for all text pairs, thereby identifying potential positive and negative samples. Subsequently, the student model is trained to ensure that the similarity between positive pairs exceeds that between negative pairs, with the $margin$ hyperparameter controlling the degree of this difference. If the batch size is $m$, the total number of text pairs (i.e., $N$) is given by $C_{C_{m}^{2}}^{2}$.
$\mathcal{L}=\lambda_{1}\mathcal{L}_{cosine}+\lambda_{2}\mathcal{L}_{sim}+\lambda_{3}\mathcal{L}_{resim}$ (4)
@@ 2.3 Stage 1&2: Distillation from Multiple Teachers
The final loss $\mathcal{L}$ is a weighted sum of the aforementioned three loss functions. where $\lambda_{1}$,$\lambda_{2}$, and $\lambda_{3}$ are hyperparameters. The biggest advantage of distillation vectors is that we do not need any supervised data. Without considering resource constraints, we can use trillions of unsupervised texts for distillation training to achieve extreme performance for a given model size.
@@
Notably, the main difference between stage 1 and stage 2 lies in the trained parameters. In stage 1, only the fully connected layer (FC1) is trained, whereas in stage 2, both the fully connected layer (FC1) and the last three encoder layers of the student model are trained.
@@ 2.4 Stage 3: Dimension Reduction
In the first two stages, the student model is trained by learning from the teacher models. Specifically, we concatenate the vectors produced by the two teacher models, resulting in a student model vector with a dimensionality of 12,288 (4,096 + 8,192), which is impractically large. Inspired by MRL Kusupati et al. (2024), we introduce three additional, independent fully connected layers (FC2, FC3, and FC4) to generate low-dimensionality vectors, each achieving a different level of dimension reduction. For instance, by incorporating the fully connected layer FC3 with a shape of (1536888This refers to the dimensionality of the encoder layer’s hidden state., 512), we obtain a more manageable 512-dimensional vector space.
For the three FC layers, since the dimensions of the reduced vectors do not align with those of the concatenated teacher vector, the $\mathcal{L}_{cosine}$ is omitted and only the $\mathcal{L}_{sim}$ and $\mathcal{L}_{resim}$ are utilized. To ensure the accuracy of the vectors generated from the FC1 layer (i.e., the 12288-dimensional vectors), they continue to be trained using all three loss functions. During this stage, all parameters of the student model are trained.
In addition to the previously mentioned dimension reduction method, we present a potentially promising approach to self-distillation, where the aligned vectors from an earlier stage of the student model’s training serve as teacher vectors. Specifically, we propose to utilize the 12288-dimensional vectors output from the FC1 layer to serve as teachers for the shorter vectors generated by the other three FC layers. This approach provides a unique advantage by enabling the reduction of the dimensionality of any embedding model, utilizing only unsupervised data and the model itself. Given that this paper primarily focuses on introducing the training methods of the Stella and Jasper models, we did not conduct experiments to evaluate the specific merits of this proposed approach.
2.5 Stage 4: Unlock Multimodal Potential
In stage 4, we leverage image-caption pairs as the training dataset, focusing exclusively on training the visual encoder while keeping the other components frozen. The training process is based on self-distillation, where the caption’s vector representation serves as the teacher vector, and the image’s vector representation acts as the student vector. All fully connected layers introduced in previous stages are employed to generate multiple pairs of student and teacher vectors. For each pair, we calculate three losses, which are then averaged to obtain the final loss.
It is important to note that this stage achieves only a preliminary alignment between the text and image modalities, leaving significant room for improvement. In future work, we aim to further explore and refine the modality alignment process.
@@ 3 Experiments
Model Model Size Average(56 datasets) Classification Clustering PairClassification Reranking Retrieval STS Summarization
NV-Embed-v2 7851M 72.31 90.37 58.46 88.67 60.65 62.65 84.31 30.7
bge-en-icl 7111M 71.67 88.95 57.89 88.14 59.86 62.16 84.24 30.77
Stella_en_1.5B_v5 1543M 71.19 87.63 57.69 88.07 61.21 61.01 84.51 31.49
SFR-Embedding-2_R 7111M 70.31 89.05 56.17 88.07 60.14 60.18 81.26 30.71
gte-Qwen2-1.5B-instruct 1776M 67.16 82.47 48.75 87.51 59.98 58.29 82.73 31.17
voyage-lite-02-instruct 1220M 67.13 79.25 52.42 86.87 58.24 56.60 85.79 31.01
Jasper (our model) 1543M+400M 71.54 88.49 58.04 88.07 60.91 61.33 84.67 31.42
Table 1: MTEB Results as of December 24, 2024. We use the original model names on the leaderboard for clarity.